# Improving Representation of High-frequency Components for Medical Foundation Models

### v2

@@ 4.1 Implementation details
We conducted experiments on two popular architectures: vanilla ViT-B [43] (86.9M parameters) and SwinT-B [19] (86.9M parameters) as the image encoders. The image decoder for reconstruction consisted of four convolutional layers, while the classification head was composed of a multilayer perception. The models were pretrained on a comprehensive dataset, including modalities of CT, MRI, X-ray, ultrasound, optical coherence tomography (OCT, projection image), retina imaging, and dermatoscopy.111The sources of the datasets are provided on the GitHub webpage.. All images were resized to using bi-cubic interpolation and zero-padding. For single-channel grayscale images, the image was duplicated along the channel dimension to maintain consistency. The model was trained using the Adam optimizer with a learning rate of . We employed image augmentation techniques including random flipping, random rotation, random affine transformations, and random histogram shifting during the pretraining phase. The models were trained on a Linux workstation with two A100 (80G) GPUs for 100 epochs.
We compared Frepa with other pretrained and task-specific foundation models, including MAE [1], MFM [2] (a self-supervised strategy performed in the frequency domain), MedSAM [5], and CLIP [3]. The backbones for MAE and MFM were ViT-B, consistent with Frepa, and were pretrained on our collected dataset. The training of CLIP followed the method proposed in [6]. For MedSAM, we used the model and checkpoint published in [5] without modification. We also included one large visual-language model, MedDr [6] (40 billion parameters) for the comparison of some classification tasks.
@@ 4.2 Experiment settings for downstream tasks
We froze the encoders and fine-tuned only the task heads for downstream tasks. To ensure a fair comparison, we employed consistent training strategies and decoder architectures across all foundation models. The loss functions used were Cross Entropy loss for classification, Dice loss for segmentation, and GIoU loss [44] for detection tasks. For classification evaluation, we utilized accuracy (ACC) and the area under the curve (AUC) metrics. Segmentation performance was assessed using the DSC. Detection ratio and IoU were used for the evaluation of detection tasks. Notably, we incorporated two modalities—OCT reconstruction data and electron microscopy (EM) data—for segmentation tasks. These modalities were not included in the pretraining phases, thereby evaluating the generalizability to real-world data. The raw datasets were randomly divided into training, validation, and testing sets in an 8:1:1 ratio. Other training settings remain the same as above.
@@ 4.4.1 Classification and detection
Tables I and II present the results of classification and lung nodule detection on 2D images and 3D volumes, respectively. It shows that Frepa, when implemented with both ViT and SwinT, overall outperforms MAE and MAF. Specifically, Frepa exhibits an average improvement on ACC of 2.3 and 2.4 over MAE, and 3.0 and 3.1 over MFM, respectively. AUC metric also shows a similar conclusion. Furthermore, when compared to the classification-specific foundation model, CLIP, Frepa demonstrates comparable performance across most tasks and excels in certain areas such as the classification of chest cancer sub-types and pneumonia sub-types. Additionally, it is noteworthy that in fine-grained tasks, such as lung nodule and lung nodule detection, Frepa achieves more substantial improvements, attributed to its superior capability of capturing high-frequency details. Specifically, For 2D images, there is a 4.0 increase in lung cancer detection ratio and a 7 improvement in IoU (Table I), while for 3D volumes, the lung nodule detection ratio improves by 3.5 and the IoU by 6.4 (Table II). Collectively, in classification tasks reliant on global-level information, Frepa achieves performance comparable to CLIP and MAE, affirming its robust global-level representation. Moreover, for tasks requiring fine-grained information, Frepa consistently outperforms, underscoring its effectiveness in capturing detailed, high-frequency information.
@@ 4.4.2 Segmentation
Tables III and IV present the segmentation results on 2D images and 3D volumes, respectively. These results encompass both large-scale anatomical segmentations, (lung, heart, and pneumonia infection in 2D segmentation, Tables III, and abdomen1K [45], AMOS22 [46], and ADCD [47] in 3D segmentation, Tables IV), as well as fine-grained and vascular-structure segmentation (OCT artery-vein [48], retinal vessels [49, 50, 51], and intercellular gaps [14] in 2D segmentation, Tables III, and pulmonary artery-vein [15] in 3D segmentation, Tables IV). In large-scale anatomical segmentations, Frepa shows a slight improvement over other models, including MedSAM, attributed to its enhanced ability to capture high-frequency details, thereby improving boundary identification. For the fine-grained segmentation tasks, Frepa exhibits more markedly superior performance, with a significant improvement ranging from +7 to +15 in DSC compared to the best-performing models. Notably, in vessel segmentation for retina and OCT (Tables III), other models yield almost failed segmentation results (Fig. 1), whereas Frepa consistently achieves exceptional results, maintaining topological structures and morphology that closely match the ground truth. Moreover, this advancement is preserved when applied to external modalities, such as the OCT reconstruction dataset and EM dataset, which exhibit entirely different imaging characteristics from other modalities. These findings robustly demonstrate Frepa’s superior representation of fine-grained information and high-frequency components, along with better generalizability than other foundation models when applied to a broader range of realistic medical tasks.
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | Frepa+ViT | Frepa+SwinT |
| CT | Lung# | DSC() | 93.241.41 | 93.092.44 | 78.5314.62 | 95.363.47 | 96.891.19 | 96.462.80 |
| Heart# | DSC() | 88.682.02 | 88.972.30 | 65.1812.52 | 91.382.32 | 91.302.70 | 90.722.59 | |
| Pneumonia infection# | DSC() | 78.309.73 | 75.4710.18 | 55.1815.52 | 71.0710.72 | 81.478.19 | 81.519.35 | |
| OCT reconstruction | OCTA artery[48] | DSC() | 74.865.94 | 64.357.29 | 13.061.85 | 71.595.13 | 78.006.44 | 79.085.79 |
| OCT vein[48] | DSC() | 70.424.42 | 63.444.98 | 12.961.21 | 66.313.88 | 76.155.11 | 77.354.20 | |
| OCT artery[48] | DSC() | 53.708.39 | 46.458.67 | 13.101.59 | 54.517.50 | 59.456.41 | 61.886.95 | |
| OCT vein[48] | DSC() | 55.947.33 | 48.2510.40 | 13.090.96 | 53.308.11 | 60.757.36 | 61.668.52 | |
| Retina | HRF[49] | DSC() | 62.024.16 | 54.214.22 | 20.982.65 | 72.392.33 | 79.882.27 | 80.012.67 |
| DRIVE[50] | DSC() | 78.753.30 | 70.824.03 | 17.313.19 | 80.524.13 | 82.903.84 | 83.563.25 | |
| CHASEDB1[51] | DSC() | 70.965.53 | 68.672.19 | 23.302.54 | 74.841.91 | 81.152.21 | 81.522.06 | |
| EM | CREMI[14] | DSC() | 68.689.77 | 61.049.49 | 23.234.88 | 69.219.77 | 78.709.40 | 79.099.16 |
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | FrepaViT | FrepaSwinT |
| CT | Pulmonary artery[15] | DSC() | 62.5913.22 | 57.9712.66 | 20.1913.55 | 64.5711.37 | 77.579.53 | 79.5810.69 |
| Pulmonary vein[15] | DSC() | 57.5710.37 | 52.4411.73 | 18.3911.22 | 61.7111.70 | 75.6510.99 | 77.852.94 | |
| Abdomen1K-Liver[45] | DSC() | 94.002.02 | 91.661.64 | 70.8710.48 | 94.881.38 | 95.141.37 | 94.541.96 | |
| Abdomen1K-Kidney[45] | DSC() | 89.958.69 | 79.1223.29 | 26.7817.54 | 91.185.17 | 92.154.87 | 92.473.87 | |
| Abdomen1K-Spleen[45] | DSC() | 91.805.61 | 89.077.08 | 43.8211.11 | 94.901.50 | 95.361.11 | 95.501.55 | |
| MRI | AMOS22[46] | DSC() | 71.235.13 | 70.606.27 | 26.0615.26 | 73.983.18 | 76.394.34 | 77.134.72 |
| ADCD[47] | DSC() | 68.2718.66 | 59.3622.50 | 18.371.07 | 71.7418.89 | 80.5419.57 | 82.9915.03 |

### v3

@@ 4 Experiments and results
Modality | Datasets | Number of images |
DeepLesion [43] | 928,020 | |
Colonography dataset [44] | 660,207 | |
MSD dataset [45] | 147,796 | |
CQ500 [46] | 72,282 | |
CT (1,860,192) | TCIA HCC-TACE [47] | 51,887 |
UCSF-PDGM [48] | 11,523 | |
BraTS2021 [49] | 79,112 | |
Duke-Breast-Cancer-MRI [50] | 773,253 | |
ACRIN-6698 [51] | 891,786 | |
MRI (1,773,510) | Colorectal-Liver-Metastases [52] | 17,836 |
CheXpert [53] | 224,316 | |
MIMIC-CXR [54] | 371,920 | |
NIH Chest [55] | 112,120 | |
COVID-19 Radiography [56] | 21,165 | |
X-ray (730,373) | COVIDGR [57] | 852 |
Breast-Lesions-USG [58] | 522 | |
Fetal Plane [59] | 12,400 | |
Muscle fascicle length tracking [60] | 42,000 | |
TN3K [61] | 3,493 | |
EBUS [62] | 1,000 | |
Ultrasound (62,325) | Micro-Ultrasound Prostate [63] | 2,910 |
OCT 2017 [64] | 84,484 | |
OCT classification [65] | 66,075 | |
OCTDL [66] | 2064 | |
OLIVES [67] | 33,533 | |
OCT (210,167) | RetinalOCT [68] | 24,000 |
ODIR-5K [69] | 5,000 | |
Aptos2019 [70] | 3,662 | |
Retina (72,081) | Diabetic Retinopathy Detection [71] | 63,419 |
ISIC2019 [72] | 77,245 | |
Dermoscopy (98,397) | Private dataset | 21,152 |
Modality | Datasets | Downstream |
Tianchi cancer [73] | Classification | |
Lung tumor [74] | Detection | |
COPD | Classification | |
Pneumonia | Classification | |
Luna16 Lung Nodule [75] | Detection | |
Lung | Segmentation | |
Heart | Segmentation | |
Pneumonia infection [76] | Segmentation | |
Pulmonary artery [15] | Segmentation | |
Pulmonary vein [15] | Segmentation | |
Abdomen1K-Liver [77] | Segmentation | |
Abdomen1K-Kidney [77] | Segmentation | |
CT | Abdomen1K-Spleen [77] | Segmentation |
AMOS22 [78] | Segmentation | |
MRI | ACDC [79] | Segmentation |
COVID19 Radiography [56] | Classification | |
Pneumonia MNIST [12] | Classification | |
Xray | COVID19 PA Dataset [80] | Classification |
Fetal Plane [59] | Classification | |
Ultrasound | Breast MNIST [12] | Classification |
OCT 17 [64] | Classification | |
OCT | OCT MNIST [12] | Classification |
Aptos2019 [70] | Classification | |
ISIC2019 [72] | Classification | |
HRF [81] | Segmentation | |
DRIVE [82] | Segmentation | |
Retina | CHASEDB1 [83] | Segmentation |
OCTA artery [84] | Segmentation | |
OCTA vein [84] | Segmentation | |
OCT artery [84] | Segmentation | |
OCT reconstruction | OCT vein [84] | Segmentation |
EM | CREMI [14] | Segmentation |
Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT | Frepa+ConvNeXt |
ACC | 0.9410.023 | 0.9030.016 | 0.9620.013 | 0.9510.026 | 0.8870.049 | 0.9590.020 | 0.9590.022 | 0.9590.015 | ||
Tianchi Cancer | AUC | 0.9820.013 | 0.9730.010 | 0.9920.004 | 0.9860.007 | 0.9700.013 | 0.9850.012 | 0.9870.012 | 0.9850.008 | |
Detection | 0.9160.012 | 0.9050.013 | 0.6380.072 | 0.8940.014 | 0.7120.035 | 0.9540.009 | 0.9430.006 | 0.9370.006 | ||
CT | Lung tumor | IoU | 0.6010.067 | 0.6130.072 | 0.2750.148 | 0.5720.099 | 0.4550.142 | 0.6620.061 | 0.6840.086 | 0.6770.070 |
ACC | 0.9650.010 | 0.9620.008 | 0.9830.004 | 0.9700.010 | 0.9680.004 | 0.9760.010 | 0.9720.007 | 0.9740.008 | ||
COVID19 Radiography | AUC | 0.9970.001 | 0.9960.002 | 0.9980.001 | 0.9970.001 | 0.9960.001 | 0.9980.001 | 0.9980.001 | 0.9980.001 | |
ACC | 0.9750.003 | 0.9690.003 | 0.9750.002 | 0.9690.006 | 0.9620.004 | 0.9770.004 | 0.9740.002 | 0.9740.004 | ||
Pneumonia MNIST | AUC | 0.9950.002 | 0.9940.002 | 0.9960.001 | 0.9930.003 | 0.9910.003 | 0.9950.002 | 0.9940.002 | 0.9940.002 | |
ACC | 0.9400.006 | 0.9350.014 | 0.9510.006 | 0.9440.011 | 0.9400.015 | 0.9440.009 | 0.9410.007 | 0.9390.012 | ||
X-Ray | COVID19 PA Dataset | AUC | 0.9810.002 | 0.9810.007 | 0.9870.004 | 0.9820.003 | 0.9830.007 | 0.9840.003 | 0.9820.007 | 0.9800.007 |
ACC | 0.8590.008 | 0.8830.009 | 0.9030.006 | 0.8900.009 | 0.8190.012 | 0.9010.004 | 0.8960.005 | 0.9010.004 | ||
OCT17 | AUC | 0.9630.021 | 0.9700.010 | 0.9850.013 | 0.9680.024 | 0.9420.019 | 0.9890.019 | 0.9850.022 | 0.9820.015 | |
ACC | 0.8990.005 | 0.8910.004 | 0.9150.005 | 0.8820.005 | 0.8440.006 | 0.9270.004 | 0.9170.006 | 0.9120.005 | ||
OCT | OCT MNIST | AUC | 0.9630.004 | 0.9570.002 | 0.9720.002 | 0.9540.003 | 0.9390.003 | 0.9750.002 | 0.9710.002 | 0.9690.002 |
ACC | 0.9340.001 | 0.9300.003 | 0.9280.003 | 0.9240.003 | 0.8600.004 | 0.9370.002 | 0.9310.004 | 0.9350.001 | ||
Fetal Ultra | AUC | 0.9650.004 | 0.9680.004 | 0.9700.003 | 0.9640.002 | 0.9100.008 | 0.9720.004 | 0.9730.003 | 0.9720.005 | |
ACC | 0.8690.017 | 0.8430.035 | 0.8870.017 | 0.8570.029 | 0.8300.050 | 0.8730.032 | 0.8660.044 | 0.8610.011 | ||
Ultrasound | Breast MNIST | AUC | 0.8390.041 | 0.8200.054 | 0.9000.036 | 0.8570.011 | 0.8220.054 | 0.8700.027 | 0.8590.024 | 0.8590.034 |
ACC | 0.7170.023 | 0.7360.019 | 0.7070.020 | 0.7400.013 | 0.6890.019 | 0.7610.011 | 0.7640.012 | 0.7540.011 | ||
Retina | Aptos2019 | AUC | 0.8730.013 | 0.8900.009 | 0.8820.019 | 0.8860.020 | 0.8450.023 | 0.9040.007 | 0.8970.009 | 0.8990.008 |
ACC | 0.7190.005 | 0.7140.011 | 0.7120.006 | 0.7160.007 | 0.6370.002 | 0.7380.004 | 0.7420.001 | 0.7440.003 | ||
Dermatoscope | ISIC2019 | AUC | 0.9180.003 | 0.9150.004 | 0.9070.005 | 0.9140.002 | 0.8560.004 | 0.9320.005 | 0.9350.003 | 0.9320.003 |
Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT | Frepa+ConvNeXt |
ACC | 0.8530.037 | 0.8160.106 | 0.8690.024 | 0.8000.053 | 0.6560.035 | 0.8960.017 | 0.8510.076 | 0.8750.028 | ||
COPD | AUC | 0.9420.062 | 0.9260.074 | 0.9360.061 | 0.9200.044 | 0.8540.075 | 0.9610.031 | 0.9560.035 | 0.9600.046 | |
ACC | 0.8600.028 | 0.8680.054 | 0.8760.026 | 0.8880.052 | 0.8040.041 | 0.8840.030 | 0.8640.017 | 0.8880.048 | ||
Pneumonia | AUC | 0.9280.061 | 0.9170.052 | 0.9350.050 | 0.9300.074 | 0.8650.095 | 0.9380.065 | 0.9290.059 | 0.9350.063 | |
Sen | 0.8460.066 | 0.8340.076 | 0.6970.146 | 0.8750.076 | 0.8160.071 | 0.8730.087 | 0.8850.091 | 0.8890.078 | ||
CT | Luna16 | Pre | 0.8730.058 | 0.8230.059 | 0.6910.125 | 0.8930.068 | 0.8290.059 | 0.8790.057 | 0.8850.099 | 0.8770.088 |
Modality | Dataset | Metric | 3D MAE | 3DFrepa+ViT | 3DFrepa+SwinT | 3DFrepa+ConvNeXt |
ACC | 0.8720.028 | 0.8930.021 | 0.8880.024 | 0.8960.032 | ||
COPD | AUC | 0.9540.05 | 0.9670.064 | 0.9620.043 | 0.9670.056 | |
ACC | 0.8800.042 | 0.8840.036 | 0.9000.032 | 0.9000.037 | ||
Pneumonia | AUC | 0.9330.058 | 0.9320.057 | 0.9330.066 | 0.9410.062 | |
Sen | 0.8930.048 | 0.9010.042 | 0.9050.067 | 0.8910.057 | ||
CT | Luna16 | Pre | 0.8880.035 | 0.9240.031 | 0.9080.048 | 0.9110.046 |
@@ 4.1 Implementation details
4.1.1 Pretrained on 2D images
We conducted experiments on three popular architectures on 2D images: vanilla ViT-B [85] (86.9M parameters), SwinT-B [19] (86.9M parameters), and ConvNeXt [18] (feature channel=128, 87.5M parameters) as the image encoders. The image decoder for reconstruction consisted of four convolutional layers, while the classification head was composed of a multilayer perception. The models were pretrained on a comprehensive dataset, including modalities of CT, MRI, X-ray, ultrasound, optical coherence tomography (OCT, projection image), retina imaging, and dermatoscopy. Table I summarizes the datasets employed for model pretraining. For CT images, we initially normalized the Hounsfield units using typical window width and level values ([-1000, 400] for the lung window, [-160, 240] for the abdomen window, [-30, 90] for the brain window, and [-300, 1200] for the bone window) and then rescaled them to [0, 1]. For other imaging modalities, we clipped the intensity values to the range between the 0.5th and 99.5th percentiles before rescaling them to [0, 1]. All images were resized to using bi-cubic interpolation and zero-padding. For single-channel grayscale images, the image was duplicated along the channel dimension to maintain consistency. The model was trained using the Adam optimizer with a learning rate of . We employed image augmentation techniques including random flipping, random rotation, random affine transformations, and random histogram shifting during the pretraining phase. The models were trained on a Linux workstation with two A100 (80G) GPUs for 100 epochs.
We compared Frepa with other pretrained and task-specific foundation models, including MAE [1], MFM [2] (a self-supervised strategy performed in the frequency domain), MedSAM [5], and CLIP [3]. The backbones for MAE and MFM were ViT-B, consistent with Frepa, and were pretrained on our collected dataset. The training of CLIP followed the method proposed in [6]. For MedSAM, we used the model and checkpoint published in [5] without modification. We also included one large visual-language model, MedDr [6] (40 billion parameters) for the comparison of classification tasks. All encoders were used solely for image encoding, with identical decoder structures applied across models. During the downstream task training phase, the parameters of all model encoders were kept fixed, and only the decoders were trained and optimized.
4.1.2 Pretrained on 3D volumes
We also conducted model pretraining experiments directly on 3D volumes, with the method described in Section 3.3.1. Specifically, we included the same three backbones in their 3D versions: ViT (dim=1024, 32.0M parameters), SwinT (feature size=36, 35.1M parameters), and 3D ConvNeXt (MedNeXt [86], feature channel=24, 34.8M parameters). The models were pretrained on all CT and MRI volumes from our established dataset (Table I). Following standard practices for handling 3D volumes, we randomly cropped the data into sub-volume cubes to address GPU memory limitations. This cube size was consistently applied during training, validation, and testing. The input channel was set to 1, as both CT and MRI modalities consist of grayscale images. We compared the 3D-pretrained Frepa model with 3D MAE [41], which directly extends the MAE strategy to 3D data. All other experimental settings were kept consistent with those used in the 2D experiments for a fair comparison.
@@ 4.2 Experiment settings for downstream tasks
We froze the encoders and fine-tuned only the task heads for downstream tasks. To ensure a fair comparison, we employed consistent training strategies and decoder architectures across all foundation models. The loss functions used were Cross Entropy loss for classification, Dice loss for segmentation, and GIoU loss [87] for detection tasks. The datasets were randomly divided into training, validation, and testing sets in a 7:1:2 ratio. For the classification and detection tasks, we employed the five-fold training strategies to ensure robust evaluation and improve evaluation generalization. Table II summarizes the datasets used for downstream task evaluation, respectively. In the tables, ”” indicates in-house datasets, while ”” denotes external datasets not utilized during the pretraining phase.
For classification evaluation, we utilized classification accuracy (ACC) and the area under the curve (AUC) metrics. Segmentation performance was assessed using the DSC. Detection ratio and IoU were used to evaluate cancer detection tasks, while sensitivity and precision were employed for the evaluation of the Luna16 lung nodule detection. This approach aligns with the metrics used in [75] and [88], with all calculations strictly following the methods outlined in their papers. Notably, we incorporated two modalities—OCT reconstruction data and electron microscopy (EM) data—for segmentation tasks. These modalities were not included in the pretraining phases, thereby evaluating the generalizability to real-world data.
During the training phase for retinal vessel segmentation, individual datasets provide only a limited amount of training data (e.g., 45 images in HRF [81], 20 images in DRIVE [82], and 28 images in the CHASEDB [83] dataset). To address this limitation, we combined these three datasets together for training while reporting the prediction results separately for each dataset.
@@ 4.4.1 Classification and detection
Tables III, IV, and V present the results of classification and detection on 2D images and 3D volumes, respectively. It shows that Frepa, when implemented with both ViT, SwinT, or ConvNeXt, overall outperforms MAE and MAF. Specifically, Frepa exhibits an average improvement on ACC of 2.3 and 2.4 over MAE, and 3.0 and 3.1 over MFM, respectively. AUC metric also shows a similar conclusion. Furthermore, when compared to the classification-specific foundation model, CLIP, Frepa demonstrates comparable performance across most tasks and excels in certain areas such as the classification of chest cancer sub-types and pneumonia sub-types. Additionally, it is noteworthy that in fine-grained tasks, such as ISIC2019 fine-grained classification and lung tumor detection, Frepa achieves more substantial improvements, attributed to its superior capability of capturing high-frequency details. Specifically, Frepa achieves a 2.5 improvement in accuracy and a 1.7 increase in AUC on the ISIC2019 dataset, while it demonstrates a 4.0 increase in lung tumor detection ratio and a 7 improvement in IoU, as shown in Table III, compared to the state-of-the-art (SOTA) results. Collectively, in classification tasks reliant on global-level information, Frepa achieves performance comparable to CLIP and MAE, affirming its robust global-level representation. Moreover, for tasks requiring fine-grained information, Frepa consistently outperforms, underscoring its effectiveness in capturing detailed, high-frequency information.
@@ 4.4.2 Segmentation
Tables VI, VII, and VIII present the segmentation results on 2D images and 3D volumes, respectively. These results encompass both large-scale anatomical segmentations, (lung, heart, and pneumonia infection in 2D segmentation, Tables VI, and abdomen1K [77], AMOS22 [78], and ADCD [79] in 3D segmentation, Tables VII), as well as fine-grained and vascular-structure segmentation (OCT artery-vein [84], retinal vessels [81, 82, 83], and intercellular gaps [14] in 2D segmentation, Tables VI, and pulmonary artery-vein [15] in 3D segmentation, Tables VII). We also provide the segmentation results in Fig. 6 for a better-visualized comparison. In large-scale anatomical segmentations, Frepa shows a slight improvement over other models, including MedSAM, attributed to its enhanced ability to capture high-frequency details, thereby improving boundary identification. For the fine-grained segmentation tasks, Frepa exhibits more markedly superior performance, with a significant improvement ranging from +7 to +15 in DSC compared to the best-performing models. Unlike other models, which produce significantly less accurate and fragmented segmentations (Fig.1, Fig.6), Frepa achieves superior performance by capturing intricate vascular networks, preserving topological structures and fine-grained details, while the results maintain superior connectivity and closely align with the ground truth, while the results. Moreover, this advancement is preserved when applied to external modalities, such as the OCT reconstruction dataset and EM dataset, which exhibit entirely different imaging characteristics from other modalities. These findings robustly demonstrate Frepa’s superior representation of fine-grained information and high-frequency components, along with better generalizability than other foundation models when applied to a broader range of realistic medical tasks.
Modality | Dataset | Metric | MAE | MFM | MedSAM | Frepa+ViT | Frepa+SwinT | Frepa+ConvNeXt |
Lung | DSC() | 93.241.41 | 93.092.44 | 95.363.47 | 96.891.19 | 96.462.80 | 96.901.90 | |
Heart | DSC() | 88.682.02 | 88.972.30 | 91.382.32 | 91.302.70 | 90.722.59 | 91.002.80 | |
CT | Pneumonia infection | DSC() | 78.309.73 | 75.4710.18 | 71.0710.72 | 81.478.19 | 81.519.35 | 80.408.50 |
OCTA500 artery | DSC() | 74.865.94 | 64.357.29 | 71.595.13 | 78.006.44 | 79.085.79 | 79.805.90 | |
OCTA500 vein | DSC() | 70.424.42 | 63.444.98 | 66.313.88 | 76.155.11 | 77.354.20 | 77.004.10 | |
OCT500 artery | DSC() | 53.708.39 | 46.458.67 | 54.517.50 | 59.456.41 | 61.886.95 | 62.508.00 | |
OCT reconstruction | OCT500 vein | DSC() | 55.947.33 | 48.2510.40 | 53.308.11 | 60.757.36 | 61.668.52 | 60.907.00 |
HRF | DSC() | 62.024.16 | 54.214.22 | 72.392.33 | 79.882.27 | 80.012.67 | 80.502.50 | |
DRIVE | DSC() | 78.753.30 | 70.824.03 | 80.524.13 | 82.903.84 | 83.563.25 | 84.003.10 | |
Retina | CHASEDB | DSC() | 70.965.53 | 68.672.19 | 74.841.91 | 81.152.21 | 81.522.06 | 80.802.50 |
EM | CREMI | DSC() | 68.689.77 | 61.049.49 | 69.219.77 | 78.709.40 | 79.099.16 | 78.208.20 |
Modality | Dataset | Metric | MAE | MFM | MedSAM | Frepa+ViT | Frepa+SwinT | Frepa+ConvNeXt |
Pulmonary artery | DSC() | 62.5913.22 | 57.9712.66 | 64.5711.37 | 77.579.53 | 79.5810.69 | 79.0810.15 | |
Pulmonary vein | DSC() | 57.5710.37 | 52.4411.73 | 61.7111.70 | 75.6510.99 | 77.857.94 | 77.958.12 | |
Liver | DSC() | 94.002.02 | 91.661.64 | 94.881.38 | 95.141.37 | 94.541.96 | 95.102.10 | |
Kidney | DSC() | 89.958.69 | 79.1223.29 | 91.185.17 | 92.154.87 | 92.473.87 | 92.604.20 | |
CT | Spleen | DSC() | 91.805.61 | 89.077.08 | 94.901.50 | 95.361.11 | 95.501.55 | 95.551.80 |
AMOS22 | DSC() | 71.235.13 | 70.606.27 | 73.983.18 | 76.394.34 | 77.134.72 | 77.504.90 | |
MRI | ADCD | DSC() | 68.2718.66 | 59.3622.50 | 71.7418.89 | 80.5419.57 | 82.9915.03 | 82.5014.80 |
Modality | Dataset | Metric | 3D MAE | 3DFrepa+ViT | 3DFrepa+SwinT | 3DFrepa+ConvNeXt |
Pulmonary artery | DSC() | 77.2010.10 | 82.108.20 | 84.809.50 | 86.509.80 | |
Pulmonary vein | DSC() | 74.309.01 | 79.102.80 | 85.507.00 | 85.008.20 | |
Liver | DSC() | 95.501.80 | 95.501.20 | 95.051.10 | 96.101.00 | |
Kidney | DSC() | 91.802.50 | 92.703.79 | 93.203.03 | 93.803.28 | |
CT | Spleen | DSC() | 92.501.80 | 94.801.40 | 96.001.66 | 96.021.19 |
AMOS22 | DSC() | 78.547.90 | 79.706.77 | 81.876.87 | 80.505.33 | |
MRI | ADCD | DSC() | 83.863.18 | 85.359.92 | 84.087.18 | 87.818.64 |